Agenda/Minutes of the VIS Meeting on 2016/8/19
2016/8/19 13:30 - @ Room 210
Participants:Takahashi, Sato, Ishizaki, Tanaka, Shoda, Mark, Ohishi, Aso, Fabian, Zhao, Akutsu
- Task list was updated.
- No type-A meeting in this week.
- EQ stop, release sensor, dummy payload: The specification document and drawings were fixed. The bidding was announced on the 17th.
- Installation procedure: The review was held on the 10th.
- Standard filters: Cheking of 6 sets were finished and shipped to Kamioka.
- BF damper: LVDT sample is under testing. The measured noise is too large (10^-6 m/rtHz).
- Need to know the weight of the BS BF too. Mark will check where the current weight number he has came from.
- Crane issue
- Aso will set up a discussion with Uchiyama-san and Miyoki-san early next week.
- A possibility is to remove the 2F floor of the -Y side of the BS chamber and lift heavy things from the 1F there.
- Send Shoda-san where to look for the stainless steel 1st wire breakers (Mark)
Past week report
- Mark, Fabian and Hirata-san cleaned glue off the dummy BS, attached Akutsu-san's new SS mirror plates, reglued the flags and prisms, and removed the gluing hardware, so the dummy BS is ready for shipping. Tatsumi-san's refinements to the tooling (especially the new flag holders) worked very smoothly. We were more generous with the glue and much more careful with the alignment of the prisms to the barrel, both during rehearsal and the final gluing.
Mark wrote up the gluing procedure as a new document E1605505. (The original plan was to do an addendum to Tatsumi-san's PR gluing procedure JGW-E1504248, but there was too much material.)
- Hirata-san organized with Takahashi-san to get the BS payload masses shipped to Kamioka. They will go in an air-ride truck on Wed 8/17 together with a load of Type A stuff, and will arrive early on Thu 8/17.
- Hirata-san shipped the collection of Type B fasteners to Kamioka, to arrive Mon 8/15.
- Hirata-san measured another standard filter and cleaned various parts for people.
- Hirata-san worked on the SR2 Inventor assembly. He spotted a problem with the orientation: given the position of the pistons in the SR2 tank (according to the assembly from Mirapro), mating the standard Type B SR chain and the chamber in the natural way puts the optic at 90° to what it needs to be. A similar problem was found in the Type B Test at Mitaka, and it was eventually worked around, but with some difficulty. Hirata-san will investigate the best way of making a correction.
- Komori-kun, Kuwahara-kun and Akutsu-san calibrated an initial batch of 7 wide-mouthed OSEMs.
- Fabian worked on the payload assembly procedure.
- Type B Past Week Summary - week of 8/15
- Hirata-san shipped the fasteners to Kamioka, and they arrived Tuesday.
- Hirata-san shipped a large load of Type A and Type B items, which Mark and Fabian received on Thursday 8/18.
- Akutsu-san shipped the first 7 OSEMs.
- Hirata-san and Ikenoue-san did a successful function test of the BS winches with the new clamp parts for the thicker wires.
- Mark, Fabian and Zhao-kun did a major cleanup and reorganization of the BS clean booth at Kamioka, putting PR2 stuff in boxes, vacuuming and wiping, and making space.
Mark, Fabian and Zhao-kun set the crossbeams of the assembly frame to their correct positions and installed the EQ stop ring and IM. IM_Installed.jpg This is the first actual installation activity for Type B. We encounter an issue with the crane vs. the cleambooth which will need to be addressed: the slot in the ceiling for the crane has a non-removable bar at the end Crane_Issue.jpg. SR2 has the same problem, but SR3 and SRM have a removable bar.
- Type B ongoing issues
Longer cables for BF->PI and SF->PI still not received, Takahashi-san to check status.
- Longer term to-do list
Decide position of BS rack (in consultation with Miyakawa-san) - document in progress: T1605413
- Do Inventor assembly for SR2 in chamber and on assembly frame (Hirata) - ONGOING.
- Write more of payload assembly procedure (Fabian) - ONGOING
- Lower breadboard installation to be incorporated into the procedure document.
- Work on BS simulation (Fabian)
- Do BS Simulink model (Mark, Kokeyama)
- Do 2D drawings of SR2 parts (Hirata)
- Redesign SRx IM wire clamps for larger prisms, remachined RMs
- Order SRM/PRM parts (Hirata/Shoda)
- Order SRx fasteners (Hirata)
- Type-Bp modification
- Saito-san made the 1st version of the 2D drawing. Though we still have small modification, we can start cost estimate.
- F7 LVDT
- Virgo LVDT test is ongoing. The measured calibration factor and noise level is worse than expected. More tuning is required.
- Installation related issues
- BF weight is lighter than the 3D model. (CAD: 50 kg, measured: 45 kg.) Shoda has not been able to find any reason yet. Further investigation will be done in next week.
- Mirapro sent us the design of the optical bench installation jigs. We sent them some request for the update.
- The design of the assembly frame is almost fixed.
- Simulation and others
- Investigate the BF weight inconsistency (Shoda)
- Some 2D drawings not related to the modification (additional masses, OB wire hook, OB wires,,,) (Hirata)
- Assembly drawing (Shoda)
- Optical bench additional parts design
- Organize the mirror preparation
- Ask Hirose-san to remove the wire breakers from PR2/3.
- Looking for space for experiment. The optical table on that now Geo-phones are will be available.
- Calibration of OSEM is planned. Takahashi-san sent a scale to Kamioka.
- Next week:
- 3D measurement of IMMT and IFI with Kasuya-san is scheduled.
Watchdog system (Tanaka)
- Considering to change the power supply part from DC-DC converter to a series regulator.
- A wheel of the fork lift dropped in a trench. The trench was covered now.
- Installation order for Type-B and Type-Bp
BS(1-3) -> SR3(3-7) -> SR2(7-11) -> SRM (11-)
PR3 (12-3) -> PRM(3-5) -> PR2(5-7)
- We need to coordinate the BS area usage with PI assembly work and the transportation of the assembly frame.
- Aso, Mark, Takahashi will discuss it F2F.
Site work plan for the next week
- Mon: BS test hanging (Mark, Fabian, Zhao, BF weight check (Shoda)
- Tue: BS test hanging (Mark, Fabian, Zhao), BF weight check (Shoda)
- Wed: BS test hanging (Mark, Fabian, Zhao), BF weight check (Shoda)
- Thu: F2F
- Fri: F2F
- Travel (Week of 8/22):
Takahashi:25 - 27
Shoda:22 - 27
- Mark: 22 - 27
- Fabian: 22 or 23 - 27
- Okutomi: 25 - 27
- Travel (Week of 8/29):
- Mark: 29 - 9/2
- Fabian: 29 - 9/2
2016/9/2(Fri) 13:30 -
Next F2F 2016/9/9